### v1

# Can Contrastive Learning Improve Class-Imbalanced Diffusion Model?

Abstract
Training data for class-conditional image synthesis often exhibit a long-tailed distribution with limited images for tail classes. Such an imbalance causes mode collapse and reduces the diversity of synthesized images for tail classes. For class-conditional diffusion models trained on imbalanced data, we aim to improve the diversity of tail class images without compromising the fidelity and diversity of head class images. We achieve this by introducing two deceptively simple but highly effective contrastive loss functions. Firstly, we employ an unsupervised InfoNCE loss utilizing negative samples to increase the distance/dissimilarity among synthetic images, particularly for tail classes. To further enhance the diversity of tail classes, our second loss is an MSE loss that contrasts class-conditional generation with unconditional generation at large timesteps. This second loss makes the denoising process insensitive to class conditions for the initial steps, which enriches tail classes through knowledge sharing from head classes. Conditional-unconditional alignment has been shown to enhance the performance of long-tailed GAN. We are the first to adapt such alignment to diffusion models. We successfully leveraged contrastive learning for class-imbalanced diffusion models. Our contrastive learning framework is easy to implement and outperforms standard DDPM and alternative methods for class-imbalanced diffusion models across various datasets, including CIFAR10/100-LT, PlacesLT, TinyImageNetLT, and ImageNetLT.
1 Introduction
Recent advances in diffusion models ddpm ; ddim have led to breakthrough in various generation tasks such as image generation zhang2023adding ; stable_diffusion ; ruiz2023dreambooth , video generation ho2022imagen ; ho2022video , image editing couairon2022diffedit ; kawar2023imagic , 3D generation poole2022dreamfusion , etc. These diffusion-based generative models rely on large-scale datasets for training, which often follow a long-tailed distribution with dominant data for head classes and limited data for tail classes. Similar to the class-imbalanced recognition model liu2022open and the class-imbalanced generative adversarial network tan2020fairgen ; rangwani2022improving ; rangwani2023noisytwins ; utlo , diffusion model also tends to overfit on tail classes leading to mode collapse. As shown in Fig. 1, the original diffusion model generates similar or even repetitive images for tail classes. We aim to increase the diversity of tail class images while maintaining the fidelity and diversity of head class images for diffusion models.
We look into a highly related and crucial problem of long-tailed image recognition, for which contrastive learning is an effective method jiang2021self ; li2022targeted . We are curious whether contrastive learning can improve long-tailed diffusion models for image synthesis. Contrastive learning is a general learning paradigm of minimizing the distance of positive pairs while maximizing the distance of negative pairs, which has attracted lots of research interest for self-supervised representation learning chen2020simple . On the other hand, diffusion model has also emerged as an effective model for representation learning fuest2024diffusion ; baranchuk2021label . We explicitly incorporate contrastive learning regularization to the diffusion model to further enhance representation learning towards better image generation.
Figure 1: To inspect overfitting on tail classes, we visualize training images (leftmost column) and synthetic images (third to rightmost columns) that are top-10 nearst neighbors sorted by distances to the anchor training image. KNN is based on L2 distances of Inception V3 embeddings. For each example, the top row is the results of DDPM ddpm , and the bottom row shows our method. The nearest neighbors from DDPM show severe mode collapse to the training image, while our method generated more diverse images. More examples are shown in the appendix A.3.3.
While at first glance, integrating contrastive learning into diffusion models seems straightforward, identifying suitable positive and negative pairs and designing an effective loss function are far from obvious. Also, training class-conditioned diffusion models involves both conditional generation and unconditional generation at multiple timesteps, which is more complicated than a feed-forward CNN or transformer. We found the devil is in the details and designed effective contrastive losses tailored for long-tailed diffusion models.
Our first loss is an InfoNCE loss infonce with negative samples only, serving as regularization for denoising diffusion probabilistic models (DDPM). Given that mode collapse during inference manifests as generated samples being overly similar, particularly for tail classes, we introduce unsupervised contrastive learning with negative samples to maximize the distances between generated images. Unlike supervised contrastive learning khosla2020supervised , unsupervised contrastive learning does not differentiate between embedding/representation for different classes. Instead, it distinguishes between representations of images of the same class, hence increasing image diversity.
Our second loss is an MSE loss designed to contrast estimated noises from conditional and unconditional generation, which effectively minimizes the KL divergence between latent distributions for conditional and unconditional generation. While such conditional-unconditional alignment seems undesirable with less controllability, a critical aspect of our approach is that our MSE loss is weighted more for larger timesteps corresponding to the initial stage of the reverse process. Given the known overfitting in conditional generation for tail classes for both GAN and diffusion models, we are motivated by the success of conditional-unconditional alignment for long-tailed GAN utlo ; transitioncgan , which facilitates knowledge sharing between head classes and tail classes. Notably, unconditional GAN generation has been observed to achieve superior FID than conditional generation under limited data transitioncgan . We adapt conditional-unconditional alignment from GAN to diffusion models. Unlike the GAN-based method utlo aligns conditional generation and unconditional generation for low-resolution images exhibiting intra-class similarity, we leverage observed image similarity during the initial denoising steps for tail classes and head classes si2024freeu . Hence, we propose to match conditional generation with unconditional generation for large timesteps.
While contrastive learning has been proposed for diffusion models to improve adversarial robustness ouyang2023improving as guidance during inference and to find semantically meaningful directions dalva2024noiseclr , we effectively leveraged contrastive learning as regularization losses for class-imbalanced diffusion models. Our work has positive societal impact for content generation of underrepresented groups.
The main contributions of this work are as follows:
  • We propose Contrastive Learning for Diffusion Model (CLDM) with imbalanced data. Our proposed contrastive losses are easy to implement with a standard DDPM training pipeline.
  • Our first loss, InfoNCE, employs unsupervised contrastive learning loss with negative samples only, enhancing within-class diversity.
  • Our second novel MSE loss aligns unconditional generation and conditional generation for the initial steps in the denoising process, facilitating knowledge sharing between head and tail classes.
  • We improve the diversity of tail class images for conditional generation while maintaining the fidelity and diversity of head class images for multiple datasets and various resolutions, including CIFAR10LT, CIFAR100LT, PlacesLT, TinyImageNetLT, and ImageNetLT. We also improve results for unconditional generation.
2 Related Work
Class-imbalanced Image Generation Generative models such as VAE vae , GAN stylegan ; biggan , and diffusion models ddpm ; stable_diffusion generate inferior images for tail classes when trained with real-world data with a long-tailed distribution. It has attracted a lot of research interest ai2023generative ; tan2020fairgen ; utlo ; cbdm ; oclt to address this issue for various types of models. Many methods address the problem of class imbalance by augmenting training data for the tail classes. A VAE is fine-tuned on tail classes under a majority-based prior ai2023generative . It is observed that GAN tan2020fairgen can amplify biases, leading to tail classes to be barely generated during inference, highlighting that the fairness in GAN needs improvement. Recent work addresses class-imbalanced diffusion models cbdm ; diffrop ; oclt by regularization losses to align or separate the distributions of synthetic images and their corresponding latent representations across different classes. For example, CBDM cbdm loss minimizes the distance of estimated noise items between these two models, the original DDPM model and a second model trained with pseudo labels which form a uniform distribution. DiffROP diffrop attempts to combine contrastive learning with diffusion model by maximizing the distance of distributions between classes. We aim to reduce overfitting for tail classes, while DiffROP diffrop reduces overlapping between different classes. The first part of our loss, the InfoNCE loss with negative pairs only, is defined for any pair of images regardless of their classes. However, DiffROP only considers pairs of images of different classes as negative pairs without regularizing images of the same class. Conceptually, we generalize DiffROP by including extra negative pairs of images within each class, hence promoting intra-class diversity.
We are inspired by UTLO utlo for long-tailed generation with GAN, which shares the latent representations of conditional GAN with unconditional GAN and implicitly shares knowledge between the head class and tail class. The motivation with UTLO utlo is that low-resolution images from GAN are similar for head classes and tail classes. We observe a similar phenomenon for denoised images in the initial steps of the denoising process, which motivates our contrastive loss. We adapt the conditional-unconditional alignment method from GAN transitioncgan ; utlo to diffusion models.
Contrastive Learning for Representation Learning Both unsupervised contrastive learning chen2020simple ; moco and supervised contrastive learning khosla2020supervised ; li2022targeted are approaches used in representation learning to produce feature embeddings by maximizing the similarity between positive pairs and minimizing the similarity between negative pairs. Unsupervised contrastive learning does not require labels and typically augments the same data to form positive pairs (chen2020simple, ; moco, ). Supervised contrastive learning incorporates class labels and pulls together all samples from the same class while pushing apart samples from different classes. Contrastive learning is an effective method for self-supervised learning moco and long-tailed recognition jiang2021self ; li2022targeted . We effectively utilized contrastive learning for class-imbalanced diffusion models. Our work is different from the contrastive-guided diffusion process ouyang2023improving , which aims to improve adversarial robustness. Another difference is that we focus on contrastive learning during training, while contrastive-guided diffusion process ouyang2023improving is about guidance for inference. While the diffusion model originally for generation tasks becomes an emerging technique for representation learning fuest2024diffusion ; baranchuk2021label , we directly integrate contrastive learning regularization to diffusion models for better generation on imbalanced data. Notably, contrastive loss with negative samples exclusively focuses on separating dissimilar instances in the embedding space, which we leverage to address mode collapse, as detailed in Sec . 3.1. BYOL (BYOL, ) is a contrastive learning method with MSE loss for positive pairs only similar to our loss formulation in Sec. 3.2.
3 Method
Figure 2: Overview of the proposed framework of contrastive learning for diffusion model (CLDM). The encoder of the noise estimation network $e(\cdot)$ encodes an image with noise $x_{t}$ to a low-dimensional latent $h$, which is decoded to noise $\epsilon$ by the decoder $d(\cdot)$. We increase the distance of negative pairs of unconditional encoded latents for different samples by an InfoNCE loss $\mathcal{L}_{nceclr}$ with negative samples only. We form positive pairs from unconditional generation and conditional generation for the same sample $x_{t}^{i}$ and utilize an MSE loss $\mathcal{L}_{mseclr}$ to minimize their distances at initial time steps.
We propose Contrastive Learning for Diffusion Model (CLDM). Fig. 2 gives an overview of our framework. As is common in contrastive learning, our framework involes both negative and positive pairs, as outlined in Sec. 3.1 and Sec. 3.2, respectively. Sec. 3.3 summarizes our overall framework.
3.1 Negative Pairs and InfoNCE loss
We observed the synthesized images of DDPM for a tail class concentrated around the limited training images in the latent space, leading to the issue of mode collapse, as shown in the Fig. 1, and the Fig. 5 in the appendix A.3.1 with visualizing the distribution of synthetic images for the tail class. By applying the InfoNCE loss with negative samples only, we intend to increase the distance of each synthesized image from other images in the latent space, hence increasing the diversity of synthetic images in particular for tail classes.
We consider InfoNCE loss with negative samples only, where the negative samples are based on noisy images from a mini-batch. Specifically, an encoder $e(\cdot)$ encodes an image with noise $x_{t}$ to a low-dimensional latent $h$, which is decoded to noise $\epsilon$ by a decoder $d(\cdot)$ in a denoising network. Our InfoNCE loss is defined for the latents $h$ with each image in a mini-batch treated as an anchor. All other images are treated as negative samples. The anchor sample $x_{t}^{i}$ and negative samples $x_{t}^{j}$ are fed into the encoder $e(\cdot)$ to get the latents $h_{anc}^{i}=e(x_{t}^{i})$ and $h_{neg}^{j}=e(x_{t}^{j})$. Our InfoNCE contrastive loss $\mathcal{L}_{nceclr}$ is defined as follows:
$\mathcal{L}_{nceclr}=-\frac{1}{|B|}\sum_{i\in B}{\log\frac{\pi^{i}_{anc}}{\pi^{i}_{anc}+\sum_{j\in{B},j\neq i}{\pi^{j}_{neg}}}},\text{with }\pi^{i}_{anc}=\exp(\frac{h_{anc}^{i}\cdot h_{pos}^{i}}{\tau}),\pi^{j}_{neg}=\exp(\frac{h_{anc}^{i}\cdot h_{neg}^{j}}{\tau}),$ (1)
where $B$ denotes a mini-batch, and $\tau$ is a temperature for softmax which we keep $0.1$ as the default setting. Note that it is common to augment an anchor image $x_{t}^{i}$ to be a positive sample in many unsupervised contrastive learning methods. Here, we don’t apply any data augmentation and directly treat the anchor latent $h_{anc}^{i}$ itself as the positive vector in the InfoNCE loss, i.e., $h_{pos}^{i}:=h_{anc}^{i}$.
3.2 Positive Pairs and MSE loss
Conditinal-unconditinal Alignment Facilitates Knowledge Sharing for Class-imbalanced GAN We take inspiration from UTLO (utlo, ), which addresses long-tailed generation with GANs, and observes that the similarity between head class and tail class images increases at lower resolutions. To share knowledge between the head class and tail class, UTLO (utlo, ) proposes unconditional GAN objectives for lower-resolution images and conditional GAN objectives for subsequent higher-resolution images. More specifically, the lower-resolution image/feature becomes less sensitive to class labels, and the second part of the generator generates the fine details of images. Utilizing unconditional generation for lower resolution is effective in increasing the diversity and quality of tail class images. In a different setting of training GAN with limited data, Transitional-GAN (transitioncgan, ) leverages unconditional generation in a two-staged training scheme, as unconditional generation can give a better FID than conditional generation when data is limited.
Conditional-unconditinal Alignment for Class-imbalanced Diffusion Model We aim to adapt the conditional-unconditional alignment approach used in GANs (utlo, ; transitioncgan, ) to diffusion models. Our key insight is that the diffusion model denoises images recursively, and the similarity between head class images and tail class images is higher during the initial timesteps of the denoising process. Our approach is different from (utlo, ), which proposes unconditional generation for lower-resolution images. We visualize the reverse processing of unconditional generation and conditional generation with different class labels of the original DDPM, starting from the same Gaussian noise, as shown in Fig. 3 with more examples in appendix A.3.4. Images from unconditional generation and conditional generation share similar low-frequency components, e.g., the coarse shape of synthesized objects and background region, especially at the initial time steps si2024freeu .
Figure 3: The leftmost column shows synthetic images with different classes but with the same initial noise or random seed. We visualize the low-frequency component and the high-frequency component from the reverse processing. It shows that low-frequency components are similar during initial time steps for different classes with the same initial noise si2024freeu . More examples are in appendix A.3.4.
Motivated by this observation, we propose to match unconditional generation and conditional generation at the initial time steps, to enable knowledge sharing between tail classes and head classes with abundant data. Another motivation is that class conditions are not necessary for all timesteps. In fact, using adaptive guidance weight with a small weight in the initial steps can yield a better FID compared to using constant guidance weight cfgweightscheduler .
From Conditional-unconditinal Distributin Maching to MSE Loss We penalize KL divergence between the conditional distribution $p_{\theta}(x_{t-1}|x_{t}^{i},c^{i})$ and unconditional distribution $p_{\theta}(x_{t-1}|x_{t}^{i})$:
$\mathcal{L}^{i,t}_{mseclr}=D_{KL}[p_{\theta}(x_{t-1}|x_{t}^{i},c^{i})||p_{\theta}(x_{t-1}|x_{t}^{i})].$ (2)
Suppose $p_{\theta}(x_{t-1}|x_{t},c)=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t,c),\sigma_{t}^{2}I),p_{\theta}(x_{x-t}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{t}^{2}I)$, then we have:
$\mathcal{L}^{i,t}_{mseclr}=\mathbb{E}[\frac{1}{2\sigma_{t}^{2}}||\mu_{\theta}(x_{t}^{i},t,c^{i})-\mu_{\theta}(x_{t}^{i},t)||^{2}]+C,$ (3)
where C is constant, and
$\mu_{\theta}(x_{t}^{i},t,c^{i})$ $=\frac{1}{\sqrt{\alpha_{t}}}x_{t}^{i}-\frac{\beta_{t}}{\sqrt{\alpha_{t}}\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t}^{i},t,c^{i}),\quad\mu_{\theta}(x_{t}^{i},t)=\frac{1}{\sqrt{\alpha_{t}}}x_{t}^{i}-\frac{\beta_{t}}{\sqrt{\alpha_{t}}\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t}^{i},t),$ (4)
where $\alpha_{t}=1-\beta_{t},\bar{\alpha}_{t}\!\!=\!\!\prod_{i=1}^{t}(1-\beta_{i})$, $\{\beta_{t}\}_{1:T}$ is the variance schedule.
Then, $\mathcal{L}_{mseclr}$ simplifies to:
$\mathcal{L}^{i,t}_{mseclr}\propto||\epsilon_{\theta}(x_{t}^{i},t,c^{i})-\epsilon_{\theta}(x_{t}^{i},t)||^{2}.$ (5)
As shown in Fig. 2, the anchor vector $h_{anc}^{i}$ is fed into decoder $d(\cdot)$ to get unconditional noise estimation $\epsilon_{unc}^{i}:=\epsilon_{\theta}(x_{t}^{i};t,\emptyset)$. Meanwhile, the anchor image $x_{t}^{i}$ is also incorporated with the class label condition $\mathbf{c^{i}}$ fed into the network to get the conditional noise estimation $\epsilon_{cond}^{i}:=\epsilon_{\theta}(x_{t}^{i};t,\mathbf{c^{i}})$. The MSE contrastive loss is defined between the unconditional and conditional estimation:  [^1] Our MSE loss is similar to CBDM loss cbdm . We minimize the distance between conditional generation and unconditional generation in image space, while CBDM minimizes distances between conditional generations for two different conditions. Our motivation is different from CBDM, but we end up with a similar but different loss. Our work offers different interpretations of previous work and makes insightful connections.
$\mathcal{L}_{mseclr}=\frac{1}{|B|}\sum_{i\in B}{\mathbb{E}_{t,x_{0}}[\frac{t}{T}\|\epsilon_{\theta}(x_{t}^{i};t,\mathbf{c^{i}})-\epsilon_{\theta}(x_{t}^{i};t,\emptyset)\|^{2}}].$ (6)
The loss is weighted linearly by timesteps $t$, so that the initial steps with large $t$ are weighted more. In other words, we align conditional generation and unconditional generation for the initial steps.
3.3 Overall Framework
Algorithm 1 Training algorithm of CLDM.
Set $\mathcal{L}_{ddpm},\mathcal{L}_{nceclr},\mathcal{L}_{mseclr}=0$
for each image-class pair $(x_{0}^{i},c^{i})$ in this batch $B$ do
     Sample $\epsilon^{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $t\sim\mathcal{U}(\{0,1,...,T\})$
     $x^{i}_{t}=\sqrt{\bar{\alpha}_{t}}x^{i}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i}$
     $\epsilon_{unc}^{i}=\epsilon_{\theta}(x_{t}^{i};t,\emptyset),\epsilon_{cond}^{i}=\epsilon_{\theta}(x_{t}^{i};t,c^{i})$
     $h_{anc}^{i}=encoder(x_{t}^{i})$
     $\pi_{anc}=\exp(h_{anc}^{i}\cdot h_{anc}^{i}/\tau)$, Set $\pi_{neg}=0$
     for $x_{t}^{j}$ in this batch with $i\neq j$ do
         $h_{neg}^{j}=encoder(x_{t}^{j})$
         $\pi_{neg}=\pi_{neg}+\exp(h_{anc}^{i}\cdot h_{neg}^{j}/\tau)$
     end for
     $\mathcal{L}_{nceclr}=\mathcal{L}_{nceclr}+(-\log\frac{\pi_{anc}}{\pi_{anc}+\pi_{neg}})$
     if Unconditional Training of CFG then
         $\mathcal{L}_{ddpm}=\mathcal{L}_{ddpm}+\|\epsilon^{i}-\epsilon_{unc}^{i}\|^{2}$
     else if Conditional Training of CFG then
         $\mathcal{L}_{ddpm}=\mathcal{L}_{ddpm}+\|\epsilon^{i}-\epsilon_{cond}^{i}\|^{2}$
         $\mathcal{L}_{mseclr}=\mathcal{L}_{mseclr}+\frac{t}{T}\|\epsilon_{unc}^{i}-\epsilon_{cond}^{i}\|^{2}$
     end if
end for
Compute the gradient with the loss
$\mathcal{L}=\frac{1}{|B|}(\mathcal{L}_{ddpm}+\alpha\cdot\mathcal{L}_{nceclr}+\gamma\cdot\mathcal{L}_{mseclr})$
Our final loss function is the sum of the original DDPM loss and our contrastive losses:
$\mathcal{L}=\mathcal{L}_{ddpm}+\alpha\cdot\mathcal{L}_{nceclr}+\gamma\cdot\mathcal{L}_{mseclr},$ (7)
where $\mathcal{L}_{ddpm}$ is a standard mse loss for matching the estimated noise item with the ground-truth noise item, same as in DDPM ddpm $\alpha$ and $\gamma$ are the hyper-parameters for our InfoNCE loss and MSE loss.
The overall algorithm framework is shown as Algorithm 1. In classifier-free guidance (CFG), the class labels $\mathbf{c}$ are randomly dropped by a specific probability $p_{uncond}$. In our setting, we keep the $p_{uncond}=10\%$, same as (cbdm, ). The parameters of the noise estimation network $\epsilon_{\theta}$ are updated by the gradient of the final loss $\nabla_{\theta}\mathcal{L}$ as in Eq. 7.
4 Experiments
4.1 Experimental Setup
DatasetsWe report results on multiple datasets, including the long-tailed versions of CIFAR10/CIFAR100 (cifar, ), TinyImageNet (tinyimagenet, ), ImageNet (imagenet, ), and Places365 (places, ), namely CIFAR10LT/CIFAR100LT (cifarlt, ), TinyImageNetLT, ImageNetLT, and PlacesLT. These are commonly used datasets in long-tailed image generation (cbdm, ; diffrop, ; oclt, ) and recognition (cifarlt, ). Same as (cbdm, ; diffrop, ; oclt, ), we adopt the same methodology presented in (cifarlt, ) to construct long-tailed version datasets with an imbalance factor of $0.01$. We keep the original image size of $32\times 32$ for CIFAR10LT/CIFAR100LT, while resizing images to $64\times 64$ for TinyImageNetLT, ImageNetLT and PlacesLT.
Baselines We compare our proposed methods with multiple baselines, including: 1) original DDPM (ddpm, ), 2) CBDM (cbdm, ), which leverages imbalanced ground-truth label and balanced pseudo-labels for co-training, and 3) OCLT (oclt, ), which facilitates feature transfer from head classes to tail classes. We also compare to a widely used data augmentation technique, Adaptive Discriminator Adaption (ADA) (karras2020training, ) for generative models on the full DDPM 1000 sampling steps.
Evaluation MetricsWe use standard metrics to evaluate both the fidelity and diversity of generated samples, including Fréchet Inception Distance (FID) (fid, ), Inception Score (IS) (is, ), FID of tail classes (FIDtail), and Kernel Inception Distance (KID) (kid, ). FID, IS, and KID are regarded as overall quality metrics, while FIDtail are more inclined to measure quality for tail classes, which are widely used in evaluating generative models on long-tailed scenes (cbdm, ; oclt, ; utlo, ; xu2024rethinking, ). More implementation details are provided in the appendix A.1.
4.2 Quantitative Results
[…]
Table 1: Our method outperforms other baselines on all datasets. We also provide the results of DDPM trained on balanced datasets, which show the upper bound of performance. All models are measured with DDIM (ddim, ) 100 steps for conditional generation with CFG. red ‘()’ shows improvement of our method over DDPM baseline (ddpm, ). blue ‘()’ shows improvement of DDPM trained on balanced version over trained on long-tailed version.
[…]
DDPM (ddpm, ) 5.93 12.25 1.64
CBDM (cbdm, ) 5.81 10.01 1.58
OCLT (oclt, ) 6.10 11.13 1.58
CLDM (Ours) 5.56 (-0.37) 10.03 (-2.22) 1.27 (-0.37)
[…]
DDPM (ddpm, ) 6.95 17.45 1.71
CBDM (cbdm, ) 6.50 17.36 1.41
OCLT (oclt, ) 6.45 17.22 1.42
CLDM (Ours) 6.24 (-0.71) 16.35 (-1.10) 1.36 (-0.35)
[…]
DDPM (ddpm, ) 18.67 40.12 6.31
CBDM (cbdm, ) 20.90 48.07 6.55
OCLT (oclt, ) 17.72 39.67 5.61
CLDM (Ours) 17.16 (-1.51) 37.88 (-2.24) 4.89 (-1.42)
ImageNetLT DDPM (ddpm, ) 16.89 35.24 7.87
CBDM (cbdm, ) 17.38 37.22 8.04
OCLT (oclt, ) 16.45 33.82 7.04
CLDM (Ours) 16.16 (-0.73) 33.93 (-1.31) 6.60 (-1.27)
PlacesLT DDPM (ddpm, ) 13.89 23.74 5.26
CBDM (cbdm, ) 15.15 26.07 5.64
OCLT (oclt, ) 13.04 22.84 4.17
CLDM (Ours) 11.99 (-1.90) 20.84 (-2.90) 3.63 (-1.63)
Table 2: We compare our method with other baselines on CIFAR10LT/CIFAR100LT datasets, with DDPM 1000 steps for conditional generation with CFG. We also report the data augmentation method ADA karras2020training , which is orthogonal to our method.
      Method       CIFAR10LT       CIFAR100LT
[…]
      DDPM ddpm       5.81       9.36       7.09       12.64
       +ADA (karras2020training, )       -       -       6.69       12.87
      CBDM (cbdm, )       5.92       9.38       6.52       12.79
      OCLT (oclt, )       5.69       9.42       6.23       13.18
      CLDM (Ours)       5.57       9.42       5.99       13.01
Results Across Various Category Intervals To see the effectiveness of our method on tail classes, we divide the PlacesLT and TinyImageNetLT datasets into three super categories: ‘Head’, ‘Body’, and ‘Tail’, with classes sorted by the number of training images. For each dataset, the top 33% classes were allocated to the ‘Head’ category, the next 34% classes to the ‘Body’ category, and the last 33% classes to the ‘Tail’ category. The percentage $P_{category}$ of training images belonging to each category is shown in the first row of Tab. 3. Our method achieves the best FID score across all three categories on PlacesLT and TinyImageNetLT datasets. For the tail classes, which are only $\sim 3\%$, our method improves their FID from 23.74 to 20.84 on PlacesLT and 40.12 to 37.88 on TinyImageNetLT, respectively. With the proposed MSE loss $\mathcal{L}_{mseclr}$ for positive pairs, our method makes the denoising process insensitive to class conditions for the initial time steps. Thus, tail classes benefit from the rich data of head classes through knowledge sharing, leading to better diversity and fidelity. Additionally, both head and body classes show improved FID scores with our method, as demonstrated in Tab. 3.
Table 3: FID score for three ‘super-categories’: ‘Head’, ‘Body’, and ‘Tail’. All models are measured with DDIM (ddim, ) 100 sampling steps for conditional generation with CFG.
[…]
PlacesLT DDPM ddpm 19.26 20.31 23.74 13.89
CBDM cbdm 21.26 21.68 26.07 15.15
OCLT oclt 19.16 19.43 22.84 13.04
CLDM (Ours) 18.15 19.27 20.84 11.99
TinyImageNetLT DDPM ddpm 21.27 33.25 40.12 18.67
CBDM cbdm 24.12 35.02 48.07 20.90
OCLT oclt 20.64 30.62 39.67 17.72
CLDM (Ours) 20.39 31.02 37.88 17.16
Table 4: Unconditional generation w/ DDIM 100 steps.
Method CIFAR10LT CIFAR100LT
[…]
DDPM 27.52 6.65 18.53 8.68
CBDM 25.60 6.70 17.06 9.00
OCLT 31.38 6.34 18.97 8.73
Ours 24.16 6.80 15.97 9.23
Class-imbalanced Unconditional Generation We also evaluate all the models for class-imbalanced unconditional generation. As shown in Tab. 4, the proposed method reduces FID from 27.52 to 24.16 for CIFAR10LT and from 18.53 to 15.97 for CIFAR100LT. Such an improvement in unconditional generation highlights the effectiveness of the proposed contrastive learning loss, particularly the unsupervised InfoNCE loss $\mathcal{L}_{nceclr}$.
4.3 Qualitative Results
(a) Synthetic images from DDPM (ddpm, ). Red ellipses show similar images generated.
(b) Synthetic images from CLDM (ours) have more diversity and are less repetitive.
Figure 4: The synthesized results for CIFAR100LT tail classes ‘rose’ and ‘table’ are shown for our method and baseline methods. All methods initiate reverse processing from the same Gaussian noise for images at corresponding grid cells. Red dashed ellipses highlight mode collapse issues observed in DDPM. Overall, our method demonstrates higher diversity and fidelity compared to DDPM baseline.
Fig. 4 shows synthetic images of baseline methods and our method for CIFAR100LT tail classes ‘rose’ and ‘table’. All methods start the denoising process from the same Gaussian noise at corresponding grid cells. Red dashed ellipses highlight the mode collapse issues observed in DDPM. With the same initial noise, our method gives synthetic images with higher diversity and fidelity compared to DDPM. For example, DDPM always generates rose images containing only one rose. Our method can generate an image containing two roses (see the sixth column of the last row). To further clarify, in the ‘table’ class, DDPM repeatedly produces near-identical images as highlighted in red ellipses. In contrast, our method produces a more varied set of table images. This suggests that our method can better capture greater diversity in image generation when compared to DDPM (ddpm, ).
Ablation Study
Table 5: Loss Function Analysis on TinyImageNetLT, sampling with DDIM 100 steps.
[…]
$\mathcal{L}_{ddpm}$ 18.67
$\mathcal{L}_{ddpm}+\alpha\cdot\mathcal{L}_{nceclr}$ 18.05
$\mathcal{L}_{ddpm}+\gamma\cdot\mathcal{L}_{mseclr}$ 17.84
$\mathcal{L}_{ddpm}+\alpha\cdot\mathcal{L}_{nceclr}+\gamma\cdot\mathcal{L}_{mseclr}$ 17.16
Table 5 shows an ablation study of the proposed losses including $\mathcal{L}_{nceclr}$ and $\mathcal{L}_{mseclr}$. Our method achieves lower FID when applying only the $\mathcal{L}_{nceclr}$ loss or only $\mathcal{L}_{mseclr}$ loss compared to the original DDPM. The best result is achieved with both losses.
Limitation of our method Our MSE loss involves both conditional generation and unconditional generation, which requires two passes of the denosing network during training and increases training time. In practice, our training time for one epoch is 1.6x of that of DDPM. However, our method doesn’t increase inference latency.
5 Conclusion
Real-world data for training image generation models often exhibit long-tailed distributions. Similar to class-imbalanced GANs (utlo, ), class-imbalanced diffusion models tend to overfit on tail classes due to data scarcity. Inspired by the success and impact of contrastive learning for long-tailed recognition, we are interested in whether contrastive learning can improve the class-imbalanced diffusion model. With well-motivated losses and carefully crafted positive/negative pairs, we effectively leveraged contrastive learning with noticeable improvement. Our framework involves two contrastive losses. Firstly, we employ an unsupervised InfoNCE loss with negative samples only to contrast the latents of different synthetic images at every denoising step, promoting intra-class diversity in particular for tail classes. Our second MSE loss contrasts class-conditional generation with unconditional generation for large timesteps. This encourages the initial denoising steps to be class-agnostic, thereby enriching tail classes through knowledge sharing from head classes–a principle demonstrated to enhance long-tailed GAN performance transitioncgan ; utlo , which we successfully adapt to diffusion models. With the two losses, our contrastive learning method outperforms standard DDPM and other state-of-the-art methods, including CBDM cbdm and OCLT oclt for class-imbalanced diffusion models. Extensive experiments on multiple datasets, including CIFAR10LT, CIFAR100LT, PlacesLT, TinyImageNetLT, and large-scale ImageNetLT, demonstrate the effectiveness and impact of our method.
Appendix A Technical Appendices and Supplementary Material
A.1 Implementation Details
Training Details We keep the same U-Net backbone used in [15] for all methods and strictly follow the training configurations of the baseline models [15, 32, 48], by setting diffusion scheduler $\beta_{1}=10^{-4}$ and $\beta_{T}=0.02$ with T=1000. Adam [49] optimizer is used to optimize the network while the learning rate is set to 0.0002 after 5000 iterations of the warmup process. We use one RTX 4090 GPU to train each model on CIFAR10LT/CIFAR100LT datasets with batch size $64$ while using 2 A100 GPUs for PlacesLT, TinyImageNetLT, and ImageNetLT datasets with batch size $128$. During the training process, all methods are trained from scratch for 300k iterations on ImageNetLT dataset, 200k iterations on CIFAR10LT/CIFAR100LT datasets, 100k iterations on TinyImageNetLT and PlacesLT datasets, while they follow the classifier-free guidance (CFG) algorithm [16], which randomly drops labels with a probability of 10%. For PlacesLT, TinyImageNetLT, and ImageNetLT datasets, we also apply timestep adaptive weight $t/T$ to infoNCE loss since we find the infoNCE loss could also benefit from such an adaptive weight, like MSE loss.
Evaluation Details To evaluate the generation performance for tail classes, we measure FID score for synthetic images of the last 30% classes in each dataset. Specifically, we categorize the ‘Tail’ classes as the last 3 classes for CIFAR10LT, the last 33 classes for CIFAR100LT, the last 66 classes for TinyImageLT (200 classes), the last 121 classes for PlacesLT (365 classes), and the last 333 classes for ImageNetLT (1000 classes), respectively. For CIFAR10LT/CIFAR100LT datasets, the evaluation metrics are based on 50k synthetic images generated by each method, while for the ImageNetLT dataset, the evaluation metrics are based on 20k synthetic images generated by each method. For TinyImageNetLT and PlacesLT dataset, the evaluation metrics are based on 10k synthetic images generated by each method. During inference, we perform a grid search algorithm for each method to determine the optimal guidance strength $\omega$ of CFG, ensuring each model achieves its best performance.
A.2 More Quantitative Results
Table 6: We compare our method with other baselines on CIFAR10LT/CIFAR100LT datasets with DDIM 10 sampling steps for conditional generation with CFG. red ‘()’ shows improvement of our method over DDPM baseline [15]. blue ‘()’ shows improvement of DDPM trained on balanced version over trained on long-tailed version.
[…]
DDPM 14.72 18.27 7.04
CBDM 13.54 16.90 6.52
OCLT 15.48 20.73 7.39
CLDM (Ours) 13.16 (-1.56) 16.83 (-1.44) 6.04 (-1.00)
[…]
DDPM 14.09 25.07 6.13
CBDM 13.37 23.97 5.83
OCLT 13.70 24.48 5.73
CLDM (Ours) 12.90 (-1.19) 23.17 (-1.90) 5.63 (-0.50)
Table 7: Statistical Analysis for random seed on CIFAR100LT, sampling with DDIM 100 steps.
[…]
0 DDPM 7.02 Ours 6.26
42 DDPM 6.95 Ours 6.24
2025 DDPM 7.04 Ours 6.27
[…]
Statistical Significance Analysis To inspect the robustness of the proposed method, we conduct statistical significance analysis on the ddpm and our method, as shown in Tab. 7. With applying different random seed, the FID deviation of DDPM is about 0.04 while our method is about 0.01, illustrating the robustness of our method for different random seeds.
A.3 More Qualitative Results
(a) Images in low-dimensional embeddings.
(b) Density distribution of synthetic images.
Figure 5: (a) Visualization of low-dimensional embeddings of five training images x for a tail class (‘worm’ in CIFAR100-LT) and synthetic images generated by DDPM [15] and our method. (b) We also show the distributions of real images and synthetic images generated by our method and the original DDPM.
A.3.1 Distribution of Synthetic Images for Tail Class
We visualize synthetic images in the feature space and their density function for our method and the original DDPM. Specifically, we use the Inception-V3 [41] model to extract $2048$ dimensional features of each image, and then apply t-SNE [43] to project these features into $2$ dimensions. As shown in Fig. 5, the original DDPM overfits and generates highly similar images, while synthetic images based on our method have more diversity. We visualize the image distribution more specifically by using kernel density estimation in the bottom. The grey region represents the distribution of real ‘worm’ images from the class-balanced CIFAR100 dataset, while red ‘x’ points denote all 5 ‘worm’ images from the class-imbalanced CIFAR100LT dataset. The red region represents the distribution of images synthesized by our method, while the blue region is for the original DDPM. Areas with higher color saturation (dark, blue, or red) indicate regions of higher density, which correspond to modes of distribution. Synthetic images from DDPM mostly concentrates around the training images, leading to mode collapse. Synthetic images from our method spans the space enclosed by all training images, the distribution of which is shown in red in (b).
Figure 6: Synthesized images of CIFAR10LT tail classes ‘horse,’ ‘ship,’ and ‘truck’ from DDPM. For each column, the images are generated starting from the same Gaussian noise. The synthesized images conditioned on different class labels exhibit similar low-frequency components.
A.3.2 Similarity of Images for Different Classes
We visualize the synthesized images of DDPM for CIFAR10LT tail classes ‘horse’, ‘ship’, and ‘truck’, as shown in Fig. 6. The synthesized images for each column in (a) and (b) are conditioned on different class labels but the same Gaussian noise. It is evident that the synthesized images of each column are coarsely similar and exhibit similar low-frequency components. For instance, in the tenth column of Fig. 6 (a), the synthesized images share a consistent background featuring a purple sky. Similarly, in the fourteenth column (the third column from right to left) of Fig. 6 (b), the images consistently depict an autumn scene with an orange sky. Such observation motivates us to develop knowledge sharing between head and tail classes inspired by the knowledge sharing for long-tailed GAN [22].
Figure 7: To see overfitting on tail classes, we find Top-10 nearest neighbors (3rd to 12th columns sorted by distances) among 1000 synthetic images to an anchor image (Leftmost column) in the training set. KNN is based on $L_{2}$ distances of Inception V3 embeddings. For each example, the top row is the results of DDPM [15], and the bottom row shows ours. The nearest neighbors from DDPM show the mode collapse to the training image, while our method generated more diverse images.
A.3.3 More Visualization of Synthetic Images
To further illustrate how the proposed method mitigates the issue of overfitting, we visualize the top-10 nearest neighbors among 1000 synthetic images to an anchor image in the training set for the original DDPM and our method. As shown in Fig. 7, DDPM shows the mode collapse to the training image while our method’s generated images show much better diversity. For example, in the first two rows, DDPM generates repeated vertical worms while our method generates worms with diverse directions. In the 9th-10th rows, DDPM generates tables with repeated modes while our method generates tables with different styles and colors. We also visualize more synthetic images of CIFAT100LT tail classes for DDPM and our method, as shown in Fig. 8. All these examples illustrate the effectiveness of our method on tail class generation with increased diversity.
(a) DDPM
(b) CLDM (ours)
Figure 8: More synthetic results for CIFAR100LT tail classes. Images in corresponding grid cell for DDPM and our method are initialized from the same Gaussian noise. Our method achieves more diverse images for tail classes.
A.3.4 More Visualization of Reverse Process of DDPM for Different Classes
To illustrate our observation in Section 3.2 more clearly, we decompose $x_{t}$ into a combination of low-frequency images and high-frequency images, as shown in Fig 10, 11. The low-frequency images with the same initial noise are very similar for different classes for the initial steps, which is also observed in prior work [39].
A.4 Dimensional Collapse in Latent Space
Figure 9: Singular values of the covariance matrix for latent embeddings of synthetic images generated by DDPM and our method. ‘t’ denotes the timestep. Compared to DDPM, our method’s gives a covariance matrix with more significant singular values, which the latent embeddings effectively span more dimensions in the available space. On the other hand, DDPM has more severe dimensional collapse in the latent space, which leads to overfitting.
Dimensional collapseis the phenomenon in representation learning when embeddings span a lower-dimensional subspace instead of the entire available space. It is considered undesirable and is highly correlated with overfitting. To investigate the effectiveness of the proposed infoNCE loss, we visualize the singular values of covariance matrix of embedding vectors for synthetic images generated by DDPM and our method, as shown in Fig. 9. We evaluate dimensional collapse by first collecting the latent vectors $h$ on 1024 synthetic images. Each latent vector has a dimension of $d=256$. Then, we compute the covariance matrix $C\in\mathbb{R}^{d\times d}$ of the latent space: $C=\frac{1}{N}\sum^{N}_{i=1}{(h_{i}-\bar{h})(h_{i}-\bar{h})^{T}}$, where $\bar{h}=\sum^{N}_{i=1}{h_{i}}/N$ and $N$ is the number of samples. Fig. 9 shows the sorted sigular values of this matrix ($C=USV^{T},S=diag(\delta^{k})$) in logarithmic scale ($\log(\delta^{k})$). For DDPM, a number of log singular values less than zero, indicating collapsed dimensions. Compared to DDPM, our method has more effective features with the log of sigular values greater than zero in the latent space with less dimensional collapse, implying higher diversity of the synthetic images.
Figure 10: Top: reverse process starting from the same initial Gaussian noise but with different class conditions (2nd-4th rows) or without condition (1st row). Middle: low-frequency components of each noisy image. Bottom: high-frequency components of each noisy image. We conclude that low-frequency components are similar for different classes with the same initial noise, particularly for initial time steps. This phenomenon is also observed in prior work [39].
Figure 11: Top: reverse process starting from the same initial Gaussian noise but with different class conditions (2nd-4th rows) or without condition (1st row). Middle: low-frequency components of each noisy image. Bottom: high-frequency components of each noisy image.